OpenStudy (anonymous):

An elastic band is hung on a hook and a mass is hung on the lower end of the band. When the mass is pulled downward and released, it vibrates vertically. The equation of motion is s=2cos t + 3 sin t, t>= 0, where s is measured in centimeters and t in seconds (The positive direction is downward). A) When does the mass pass through the equilibrium position for the first time? (Equilibrium is the starting position before being pulled downward). .. A way to do this without using parametrics?

OpenStudy (anonymous):

at equilibrium position, s=0 so, 2cos t+3sin t=0 hence, tan t=-2/3 find out the least positive value of t

OpenStudy (anonymous):

When does tan t = -2/3? How do I solve that?

OpenStudy (anonymous):

find out \[\tan^{-1} (-2/3)\] by using a calculator in radian form. u will get a negative result. add pi to it. it will be ur answer

OpenStudy (anonymous):

Why would I have to add pi?